OpenStudy (anonymous):

A circle has its center at (-1, 2) and a radius of 3 units. What is the equation of the circle? (x - 1)2 + (y + 2)2 = 3 (x + 1)2 + (y - 2)2 = 3 (x + 1)2 + (y + 2)2 = 9 (x + 1)2 + (y - 2)2 = 9

OpenStudy (lgbasallote):

hint: i'll give you a demo a circle has center (-3, 4) and radius 2 equation of circle is \[(x-h)^2 + (y-k)^2 = r^2\] where (h,k) is the center and r is the radius since in my example h is -3 and k is 4 i will have \[(x- - 3)^2 + (y-4)^2 = 2^2\] \[\implies (x +3)^2 + (y-4)^2 = 4\] do you get the idea?

OpenStudy (anonymous):

Standard equation is given by: \[\large (x-h)^2 + (y-k)^2 = r^2\] h. k are the centre and r is the radius and try to find it..

OpenStudy (anonymous):

idk

OpenStudy (lgbasallote):

what is the coordinate of your center in this problem?

OpenStudy (anonymous):

i think its D ?

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

thanks....

OpenStudy (anonymous):

But how ???

OpenStudy (anonymous):

i followed your example

OpenStudy (lgbasallote):

\[(x- -1)^2 + (y-3) = 3^2\] \[\implies (x+1)^2 + (y-3) = 9 \checkmark\] you're right

OpenStudy (anonymous):

thanks(: